### v1

# Exploring Scene Coherence for Semi-Supervised 3D Semantic Segmentation

Abstract
Semi-supervised semantic segmentation, which efficiently addresses the limitation of acquiring dense annotations, is essential for 3D scene understanding. Most methods leverage the teacher model to generate pseudo labels, and then guide the learning of the student model on unlabeled scenes. However, they focus only on points with pseudo labels while directly overlooking points without pseudo labels, namely intra-scene inconsistency, leading to semantic ambiguity. Moreover, inter-scene correlation between labeled and unlabeled scenes contribute to transferring rich annotation information, yet this has not been explored for the semi-supervised tasks. To address these two problems, we propose to explore scene coherence for semi-supervised 3D semantic segmentation, dubbed CoScene. Inspired by the unstructured and unordered nature of the point clouds, our CoScene adopts the straightforward point erasure strategy to ensure the intra-scene consistency. Moreover, patch-based data augmentation is proposed to enhance the inter-scene information transfer between labeled and unlabeled scenes at both scene and instance levels. Extensive experimental results on SemanticKITTI and nuScenes show that our approach outperforms existing methods.
Introduction
LiDAR (Light Detection and Ranging) point cloud semantic segmentation, aiming to recognize the class label for each point of an input point cloud scene, is a core task in autonomous driving. Due to the advancement of deep neural networks, tremendous progress has been achieved (Hu et al. 2020; Zhu et al. 2021; Su et al. 2023). However, the excellent performance of most studies depends heavily on large-scale datasets with high-quality point-level annotations. Such dense and delicate point-level labeling is labor-intensive and time-consuming, hindering the deployment of point cloud semantic segmentation models in real-world applications. Therefore, there is a compelling need to reduce the dependence on huge exhaustively labeled datasets.
Figure 1: (a) Generated pseudo-label scenes from unlabeled scenes containing points with and without pseudo labels simultaneously (green and gray). (b) Regardless of whether point-based or voxel-based segmentation networks are employed, resulting in inconsistencies between forward and backward propagation. (c) Annotation information transfer between labeled and unlabeled scenes.
Semi-supervised learning (SSL), a significant machine learning paradigm for alleviating data annotation demands, seeks to utilize a small number of labeled samples alongside a vast collection of unlabeled samples to optimize deep models. Recently, several semi-supervised 3D semantic segmentation methods (Kong et al. 2023b; Jiang et al. 2021; Li, Shum, and Breckon 2023) have been proposed, with a core focus on effectively transferring diverse semantic information. These methods typically first obtain pseudo-labeled scenes from unlabeled data using the teacher network, and then feed the entire pseudo-labeled scenes into the student network for training. Although these methods have reported promising results, two main issues still need to be addressed.
As shown in Figure 1 (a), the pseudo-labeled scene consists of points with and without pseudo-labels simultaneously. During training, all points are involved in the model’s forward propagation, while only points with pseudo-labels are used for backward propagation, and other points are overlooked. Inconsistency between forward and backward propagation within the scene leads to semantic ambiguity, meaning the segmentation model is unable to determine the semantic labels for points without pseudo-labels. Consequently, this limits the effectiveness of semantic information transfer. Both point-based (Hu et al. 2020) and voxel-based (Choy, Gwak, and Savarese 2019; Zhu et al. 2021) methods suffer from this issue, as illustrated in Figure 1 (b).
Diverse semantic information is crucial for the success of semi-supervised learning of the segmentation models (Kong et al. 2023b; Xiao et al. 2022). Thus, many studies focus on developing data augmentation methods (Sheshappanavar, Singh, and Kambhamettu 2021; Zhang et al. 2022) to expand the data distribution. However, these methods typically model the correlation between two scenes, while neglecting the correlation among multiple scenes. As shown in Figure 1 (c), the labeled and unlabeled scenes may exhibit similarities in factors such as spatial structures. Exploring the correlations between scenes can be beneficial for annotation information transfer, thereby enhancing learning on unlabeled scenes.
To address the above two issues, we explore Scene Coherence and propose CoScene for semi-supervised 3D semantic segmentation. Specifically, based on the unstructured and unordered nature of point clouds, we design a point erasure strategy that removes the points without pseudo labels during forward propagation, achieving full annotation of the training scene and ensuring intra-scene consistency. This strategy complements the insufficiency of the basic pseudo-label mechanism, which can serve as a plug-and-play module for semi-supervised 3D segmentation frameworks based on pseudo labels. Additionally, we design a patch-based data augmentation to model the correlation between multiple scenes at both the scene and instance levels. The key idea is to construct two patch pools to store scene-level and instance-level information, and then combine scene-level and instance-level patches from multiple scenes to generate diverse training samples, facilitating the transfer of annotation information to unlabeled scenes.
The proposed CoScene has been extensively evaluated through experiments and has demonstrated superiority across various benchmarks for semi-supervised 3D semantic segmentation. Under the setting of 5% labeled scenes, our method outperforms the state-of-the-art method by 4% in terms of mIoU on the SemanticKITTI (Behley et al. 2019) dataset. To summarize, our contributions are as follows:
  • We propose a plug-and-play point erasure strategy complementing with pseudo-label mechanism in semi-supervised point cloud semantic segmentation, ensuring intra-scene consistency.
  • We propose a patch-based data augmentation to model inter-scene correlation information at scene and instance levels simultaneously, facilitating information transfer between labeled scenes and unlabeled scene.
  • Extensive experiments demonstrate that our semi-supervised 3D segmentation methods achieves state-of-the-art performance, and comprehensive ablations studies are conducted to demonstrate its effectiveness.
Figure 2: Overview of the proposed CoScene training pipeline. The teacher network generates the pseudo-labels $y_{i}^{u}$ from the unlabeled scene $x_{i}^{u}$, followed by an erasure operation to filter out low-confidence points, ensuring training consistency. For the labeled scene $x_{i}^{l}$ and the generated erased scene $\hat{x}_{i}^{u}$, we employ a patch-based data augmentation to obtain augmented scenes $x_{i}^{lpf}$ and $\hat{x}_{i}^{upf}$, separately. Subsequently, both scenes are fed into the student network for computing the loss of $\mathcal{L}_{l}$ and $\mathcal{L}_{u}$.
Related Work
LiDAR Semantic Segmentation
LiDAR semantic segmentation aims to recognize the class label for each point of an input point cloud scene. Through the development of various deep neural networks, numerous approaches for LiDAR semantic segmentation have been explored over the past few years, which can be broadly grouped into three categories, i.e., voxel-based (Zhu et al. 2021), 2D projection-based (Kong et al. 2023a; Zhao, Bai, and Huang 2021), and point-based methods (Hu et al. 2020; Qi et al. 2017). Prior fully supervised works have made significant progress and achieved appealing performance, while the results deeply depend on large-scale precisely annotated datasets (Behley et al. 2019; Fong et al. 2022). To reduce annotations of 3D data, some methods with less supervision have been explored recently, e.g., weakly-supervised (Pan et al. 2024; Liu et al. 2022; Unal, Dai, and Van Gool 2022; Zhang et al. 2021; Chen et al. 2023; Tang et al. 2024) and semi-supervised learning (Li, Shum, and Breckon 2023; Reichardt, Ebert, and Wasenmüller 2023; Zou et al. 2018; Chen et al. 2021). LaserMix (Kong et al. 2023b), as the pioneer in semi-supervised LiDAR segmentation leverages the simple laser beam mix strategy to get solid results. Follow-up work fuses LiDAR-Camera data to mine deeper information (Chen et al. 2024). For semi-supervised approaches, the overall architecture often builds upon the mean-teacher paradigm (Tarvainen and Valpola 2017) and leverages the concept of pseudo label (Arazo et al. 2020). However, many points without pseudo labels contribute equally to pseudo-label points in the forward propagation, but are directly ignored when performing backward propagation, which brings intra-scene inconsistency issues. Hence, our CoScene leverages the point erasure strategy to mitigate this inconsistency problem, creating a complement to the pseudo-label mechanism to enable effective information transfer.
3D Point Cloud Data Augmentation
Point cloud data augmentation has been explored across various 3D vision tasks (Zhu, Fan, and Weng 2023; Shin et al. 2023; Wang et al. 2024; Fang et al. 2021). Enlightened by the success of Mixup (Zhang et al. 2017) in image classification, several works explore the techniques for mixing selected point cloud samples. PointMixup (Chen et al. 2020) leverages interpolation to generate mixed samples for training. To preserve structural information, RSMix (Lee et al. 2021) provides a neighboring function, and PointCutMix (Zhang et al. 2022) establishes the correspondence relationship for sample fusion. Notably, these works aim at local object-level augmentation. As a result, several works also study the mixing strategies for point cloud scene augmentation. Mix3D (Nekrasov et al. 2021) concatenates points of two scenes to create a new training sample. PloarMix (Xiao et al. 2022) swaps points within the fan-shaped area between two scenes. CoSMix (Saltori et al. 2023) aims to mitigate the domain shift by mixing semantic point groups. These works only support similarity modeling between two scenes, but fail to construct correlation among multiple scenes. By contrast, our approach has the capability to model inter-scene information from multiple scenes to create more diversified data, while also extracting instance-level semantic information for individual classes to further enhance scenes. Therefore, our approach can effectively guide the transfer of annotation information from labeled scenes to unlabeled scenes, facilitating semi-supervised learning.
Method
In this section, we first state the definition of semi-supervised 3D semantic segmentation and introduce the mainstream teacher-student framework. Then, we present the intra-scene consistency and inter-scene correlation of CoScene, along with its core component, i.e., the point erasure strategy and patch-based data augmentation. Finally, we outline the overall training objective.
Preliminary
Problem Formulation. Before we describe our model and training process, we first provide the definition of semi-supervised point cloud segmentation. In detail, we have a small labeled training set $D_{l}=\left\{x_{i}^{l},y_{i}^{l}\right\}_{i=1}^{N_{l}}$ and a large unlabeled training set $D_{u}=\left\{x_{i}^{u}\right\}_{i=1}^{N_{u}}$, where $x_{i}\in\mathbb{R}^{m\times{(3+r)}}$ is a point cloud scene of $m$ points with three-dimensional Cartesian coordinates relative to the sensor and additional $r$-dimensional feature (e.g., intensity, color, or reflectivity), $y_{i}$ represents its corresponding point-level segmentation labels, $N_{l}$ and $N_{u}$ are the numbers of labeled and unlabeled point cloud scenes, respectively, and $N_{u}\gg N_{l}$ in ideal cases. The goal of SSL-based point cloud segmentation is to learn a segmentation model using the limited labeled data and large-scale unlabeled data.
Teacher-Student Paradigm. Following current mainstream research (Li, Shum, and Breckon 2023), our learning framework also builds upon the teacher-student paradigm, which includes two 3D segmentation networks with the same architecture. Here, we adopt Cylinder3D (Zhu et al. 2021) as the segmentation backbone $\Phi(\cdot)$ to enable a fair comparison with other methods. In detail, the teacher model generates pseudo-labels for unlabeled scenes and undergoes a gradual update through the exponential moving average (EMA) (Tarvainen and Valpola 2017) of the student weights:
$\theta_{\mathbb{t}}^{k+1}=\alpha\cdot\theta_{\mathbb{t}}^{k}+(1-\alpha)\cdot\theta_{\mathbb{s}}^{k},$ (1)
where $\alpha$ is the EMA decay rate, which is set to 0.99 by default, $\theta_{\mathbb{t}}$ and $\theta_{\mathbb{s}}$ represent the weight parameters of the teacher and student models, respectively, and $k$ denotes the training step.
Point Erasure Strategy
Based on the teacher-student mutual learning framework (Sohn et al. 2020), the overall architecture of CoScene is depicted in Figure 2. Specifically, CoScene consists of two primary branches: one dedicated to labeled scenes and the other to unlabeled scenes. How to effectively leverage a large number of unlabeled scenes is crucial for semi-supervised learning. For the branch of unlabeled scenes, we use a teacher network to generate pseudo-labeled scenes $(x_{i}^{u},y_{i}^{u})$ based on a predefined confidence threshold $\tau_{s}$ (Kong et al. 2023b). As mentioned previously, when using the pseudo-labeled scene for training, points without pseudo labels are involved in the forward propagation but are ignored in the backward propagation due to the lack of corresponding supervision. This leads to intra-scene inconsistency, which limits the effectiveness of semantic information transfer. To address this issue, we propose erasing points without pseudo labels, resulting in an erased scene $(\hat{x}_{i}^{u},\hat{y}_{i}^{u})$ that contains fully annotated information, similar to fully supervised scenes. Formally, intra-scene inconsistency addressed by the point erasure strategy can be defined as:
$\hat{x}_{i}^{u}=\{x_{i}^{u}|\Phi_{s}(x_{i}^{u})\geq\tau_{s}\},$ (2)
where $\hat{x}_{i}^{u}$ is the erased scene with corresponding class labels $\hat{y}_{i}^{u}$, $\Phi_{s}(\cdot)$ represents the segmentation network that predicts the confidence score for each point in the scene $x_{i}^{u}$, and $\tau_{s}$ is a predefined confidence threshold, equivalent to the pseudo-label mining threshold, requiring no additional adjustment. Since the proposed point erasure strategy is independent of the segmentation network architecture, it can be utilized as a plug-and-play tool for semi-supervised 3D semantic segmentation methods based on the pseudo-label mechanism (see Figure 5 (a)).
Figure 3: Illustration of the pool initialization process. For each labeled point cloud scene, splitting it into multiple patches from the BEV perspective and assigning different instance categories to each patch enable the extraction of instances belonging to different classes. Subsequently, these patches and instances are stored individually in the patch pool and instance pool.
Patch-based Data Augmentation
Based on the labeled and erased scenes, we propose a patch-based data augmentation to enhance the transfer of annotation information, guided by inter-scene correlation. The key idea is to construct patch pools to store scene-level and instance-level information, and then combine patches from multiple scenes to generate diverse training samples.
Pool Initialization. We first use the labeled training set $D_{l}$ to create two patch pools, which store local-scene and single-instance information, respectively, as shown in Figure 3. Given the $i$-th point cloud scene $x_{i}^{l}$, we patchify $x_{i}^{l}$ into a sequence of $T$ regular, non-overlapping point patches $\{p_{ij}^{l}\}_{j=1}^{T}$, each of size $(n\times n)$, in the bird’s eye view. This process can be expressed as:
$\left\{p_{ij}^{l}\right\}_{j=1}^{T}=\text{patchify}(x_{i}^{l}),$ (3)
where $p_{ij}^{l}$ denotes the $j$-th point patch from the sequence. From all labeled scenes, we then construct the patch pool $\mathcal{B}_{\text{Pat}}^{l}$, which can be defined as:
$\mathcal{B}_{\text{Pat}}^{l}=\left\{\left\{p_{ij}^{l}\right\}_{j=1}^{T}\right\}_{i=1}^{N_{l}}.$ (4)
Based on the patch pool, we can obtain the instance pool $\mathcal{B}_{\text{Ins}}^{l}$ by assigning different class labels to each patch, which contains instance information for different categories (e.g., car and person). Following the same steps used for labeled scenes, we further create a pseudo patch pool $\mathcal{B}_{\text{Pat}}^{u}$ and a pseudo instance pool $\mathcal{B}_{\text{Ins}}^{u}$ from the erased scene $\hat{x}_{i}^{u}$. Unlike $\mathcal{B}_{\text{Pat}}^{l}$ and $\mathcal{B}_{\text{Ins}}^{l}$ covering the whole training set $D_{l}$, the pseudo pools $\mathcal{B}_{\text{Pat}}^{u}$ and $\mathcal{B}_{\text{Ins}}^{u}$ only contain scene and instance information from the current batch and are reset in each iteration. This prevents low-quality pseudo labels from persisting in pools and misleading model training. Additionally, if the number of points within a patch or instance falls below a certain threshold $\tau_{\text{min}}$, it will not be saved into the pool, ensuring effective inter-scene information transfer.
Figure 4: (a) Illustration of the MixPatch process. (b) Illustration of selecting instances from the patch pool for filling. An instance will not be filled if there are no other points around the filling location or if the filled instance overlaps with existing instances.
MixPatch. To model the inter-scene correlation and then guide the transfer of annotation information between labeled scenes and unlabeled scenes, we combine multiple scenes using patch pools, as shown in Figure 4. Figure 4 (a) illustrates the patch mixing (MixPatch), while Figure 4 (b) presents the instance filling (InsFill), which is introduced in the next paragraph. During a training iteration, the mini-batch is comprised of labeled and unlabeled scenes. After the unlabeled scene $x_{i}^{u}$ is transformed into the erased scene $\hat{x}_{i}^{u}$ using the point erasure strategy, we integrate information from the patch pool into $\hat{x}_{i}^{u}$. The patch mixing process can be described as follows:
$\hat{x}_{i}^{up}=(\hat{x}_{i}^{u}\odot\mathcal{M})\oplus(\mathcal{B}_{\text{Pat}}^{l}\odot(1-\mathcal{M})),$ (5)
where $\mathcal{M}$ is a binary mask indicating which patches from $\hat{x}_{i}^{u}$ will be replaced with selected patches of the same patch index from the patch pool, $\odot$ denotes element-wise multiplication, and $\hat{x}_{i}^{up}$ represents the augmented scene. For labeled scenes, MixPatch is applied based on the current pseudo patch pool $\mathcal{B}_{\text{Pat}}^{u}$, which can be expressed as the following:
$x_{i}^{lp}=(x_{i}^{l}\odot(1-\mathcal{M}))\oplus(\mathcal{B}_{\text{Pat}}^{u}\odot\mathcal{M}),$ (6)
where $\mathcal{M}$ is the same as defined in Eq. (5), thereby ensuring the full utilization of scene information from the unlabeled scenes. The binary mask $\mathcal{M}$ and its complement $1-\mathcal{M}$ are used to combine inter-scene point cloud patches with the erased scene, facilitating the transfer of annotation information between labeled and unlabeled scenes. To alleviate bias in the patch mixing process, we apply a uniform sampling strategy on the patches of each scene, ensuring that each patch has an equal opportunity to be selected in the mixed scene.
InsFill. In addition to patch mixing, we utilize the instances stored in the instance pool $\mathcal{B}_{\text{Ins}}^{l}$ to further enhance the scene $\hat{x}_{i}^{up}$. Given a patch with location index $j$ in the scene, we randomly select an instance from the instance pool $\mathcal{B}_{\text{Ins}}^{l}$ that was extracted from the $j$-th patch. Patches with identical indices typically correspond to similar LiDAR spatial positions, ensuring consistency between the selected instance and its fill positions in the scene. As depicted in Figure 4 (b), if the selected instance overlaps with existing instances in the scene or lacks sufficient contextual information, the filling operation is bypassed. This judgment process can be easily implemented using the maximum 3D bounding boxes. This data augmentation provides single-instance semantic information for the point cloud scene, enabling the augmented scene $\hat{x}_{i}^{upf}$ and corresponding labels $\hat{y}_{i}^{upf}$ to effectively support semi-supervised training. For the labeled branch, we also apply the same filling strategy as described, selecting instances from the pseudo instance pool $\mathcal{B}_{\text{Ins}}^{u}$ to fill the current scene. This process yields the scene $x_{i}^{lpf}$ along with its corresponding label $y_{i}^{lpf}$.
Model Training
Following the standard teacher-student mutual learning paradigm (Kong et al. 2023b; Sohn et al. 2020), we construct a mini-batch of data with an equal composition of labeled and unlabeled scenes at each iteration. The unlabeled scenes are assigned with pseudo-labels by the teacher model and then train the student model jointly with labeled scenes. We use three types of loss functions to supervise the model training here, which is shown as follows:
$\ \mathcal{L}_{s}=\mathcal{L}_{seg}(\Phi_{\theta_{s}}(x_{i}^{l}),y_{i}^{l}),$ (7)
$\ \mathcal{L}_{l}=\mathcal{L}_{seg}(\Phi_{\theta_{s}}(x_{i}^{lpf}),y_{i}^{lpf}),$ (8)
$\ \mathcal{L}_{u}=\mathcal{L}_{seg}(\Phi_{\theta_{s}}(\hat{x}_{i}^{upf}),\hat{y}_{i}^{upf}),$ (9)
where $\mathcal{L}_{s}$, $\mathcal{L}_{l}$, and $\mathcal{L}_{u}$ denote the segmentation loss for the labeled scene $x_{i}^{l}$, augmented labeled scene $x_{i}^{lpf}$, and augmented unlabeled scene $\hat{x}_{i}^{upf}$, respectively, while $\mathcal{L}_{seg}$ is the original loss of the segmentation network $\Phi(\cdot)$. We sum up all the losses together as our final training target that
$\mathcal{L}=\mathcal{L}_{s}+\lambda_{u}\mathcal{L}_{u}+\lambda_{l}\mathcal{L}_{l},$ (10)
where $\lambda_{u}$ and $\lambda_{l}$ are loss weights that balance the two loss terms. After that, we update the teacher model weights via the EMA strategy described in Eq. (1).
SemanticKITTI nuScenes
Method Venue 1% 5% 10% 20% 40% 50% 1% 10% 20% 50%
Cylinder3D CVPR 2021 45.4 45.4 56.1 57.8 58.7 58.7 50.9 65.9 66.6 71.2
MeanTeacher NeurIPS 2017 45.4 - 57.1 59.2 - 60.0 51.6 66.0 67.1 71.7
CBST ECCV 2018 48.8 - 58.3 59.4 - 59.7 53.0 66.5 69.6 71.6
CPS CVPR 2021 46.7 - 58.7 59.6 - 60.5 52.9 66.3 70.0 72.5
GPC ICCV 2021 - 41.8 49.9 58.8 59.9 - - - - -
ScribbleKITTI CVPR 2022 - 49.9 58.7 59.1 60.9 - - - - -
LiM3D CVPR 2023 58.4 59.5 62.2 63.1 63.3 63.6 - - - -
CylinderMix* AAAI 2024 51.3 56.5 60.8 62.0 62.7 62.5 55.7 69.5 72.0 73.3
LaserMix CVPR 2023 50.6 56.7 60.0 61.9 62.3 62.3 55.3 69.9 71.8 73.2
Ours (CoScene) - 54.5 62.1 63.3 63.7 63.8 64.3 56.6 70.2 72.8 73.9
Table 1: Experimental results on the SemanticKITTI and nuScenes dataset compared with different SSL methods. The results are reported with mIoU scores (%) and the best score for different split is marked in bold. “*” denotes that we reproduce the CylinderMix in the LiDAR branch (Chen et al. 2024) for fair comparison.
Experiments
Datasets and Evaluation Metrics
SemanticKITTI. Following state-of-the-art methods (Kong et al. 2023b; Chen et al. 2024), we evaluate our CoScene on the SemanticKITTI dataset (Behley et al. 2019), which is a large-scale segmentation dataset consisting of LiDAR data derived from the popular KITTI dataset (Geiger, Lenz, and Urtasun 2012). It is composed of 22 sequences, with sequences 00-10 serving as the train set (of which 08 as the val set) and sequences 11-21 as the test set. Moreover, all the point cloud scenes are collected from real-world rural and urban road environments in Germany, and the semantic class labels are mapped to 19 classes in our experiments.
nuScenes. The nuScenes dataset (Fong et al. 2022) is another well-known large autonomous driving dataset. This dataset is composed of 1000 sequences of 20s duration (850 for training and validation, 150 for testing), which are captured by the LiDAR sensor from Boston and Singapore. Following the official splitting protocol, the train and val sets contains 28,130 and 6,019 point cloud scenes, respectively, including per-points annotation for 16 semantic classes.
Evaluation Metrics. For both the SemanticKITTI and nuScenes datasets, we sample different ratios of annotated train scenes to obtain semi-supervised data and ensure the sampling indices are the same as in LaserMix(Kong et al. 2023b) for fair comparison. Meanwhile, we report the performance on the validation set using the mean Intersection over Union (mIoU) metric as the previous works (Li, Shum, and Breckon 2023; Zhu et al. 2021).
Implementation Details
We conduct experiments with 4$\times$ NVIDIA V100 GPUs with 32GB RAM, based on MMDetection3D (Contributors 2020). Inspired by the GT-sampling strategy (Yan, Mao, and Li 2018), we set $\tau_{min}$ to 5 in our experiments. For the threshold $\tau_{s}$ for pseudo-label generation and point erasure, we use the same parameter setting as in LaserMix. To facilitate patch-based data augmentation, we restrict our focus to point clouds within the X and Y dimensions in the range of $[-50,50]m$ for both the SemanticKITTI and nuScenes datasets. Similar to mainstream work, we also adopt the consistency loss proposed in MeanTeacher (Tarvainen and Valpola 2017) to align the predictions of the teacher and student networks. Additionally, the weights of the consistency loss are assigned as $1\times 10^{3}$ and $2\times 10^{3}$ for the SemanticKITTI and nuScenes datasets, respectively. The weights for $\lambda_{u}$ and $\lambda_{l}$ are set as 1 for both datasets.
Main Results
In this subsection, we compare our method with current SOTA semi-supervised point cloud segmentation methods, using sub-sampling of the datasets under various partition protocols (Kong et al. 2023b). Specifically, in the SemanticKITTI dataset (Behley et al. 2019), we split the entire training set into labeled subsets with ratios of 1%, 5%, 10%, 20%, 40%, and 50%. Similarly, in the nuScenes dataset (Fong et al. 2022), we utilize labeled subsets with ratios of 1%, 10%, 20%, and 50%. We provide detailed comparative results on both datasets against different baseline methods.
Results on SemanticKITTI Dataset. We present the comparison results between our method and state-of-the-art methods on the widely adopted SemanticKITTI dataset in Table 1. It can be clearly seen that our method effectively exploits unlabeled scenes compared to the fully supervised learning baseline Cylinder3D (Zhu et al. 2021), with a dramatic performance boost. Specifically, in the 5% ratio, our approach surpasses the fully supervised baseline by 16.7%. In other settings, our method also shows consistent improvements ranging from 5.6% to 7.2% for ratios of 10%, 20%, and 50%. In addition, comparison to the recently proposed SOTA method LaserMix (Kong et al. 2023b), which is considered a strong SSL-based point cloud segmentation method, our approach achieves improvements of 2% to 6% in all cases. The results also demonstrate that our approach outperforms the current SOTA LiM3D (Li, Shum, and Breckon 2023) in the majority of settings. In some partition protocols, our approach shows superior performance compared to the LiM3D with fewer labeled scenes. For example, our approach trained with 10% labeled scenes outperforms LiM3D trained with 40% labeled scenes. It is worth mentioning that our approach utilizes uniform sampling, similar to LaserMix, to acquire labeled scenes, without utilizing the keyframe sampling method used in LiM3D (Li, Shum, and Breckon 2023), which typically yields higher gains when the amount of labeled data is relatively low.
Results on nuScenes Dataset. As shown in Table 1, our approach still shows a significant improvement in performance compared to state-of-the-art methods on the comprehensive nuScenes dataset (Fong et al. 2022). For instance, our approach outperforms LaserMix (Kong et al. 2023b) by 1.3% and 1.0% at 1% and 20% ratios, respectively, demonstrating the generalization ability of our method.
Ablation Study
In this subsection, we study the effects of intra-scene consistency assurance through the point erasure strategy (PtErase), as well as the exploration of inter-scene correlation, incorporating patch-based mix (MixPatch) and patch-based instance filling data augmentation (InsFill) in our method. Without loss of generalizability, all experiments are conducted on the SemanticKITTI dataset with 1%, 10%, 20% and 50% ratio, and evaluated on val set. Table 2 demonstrates the improvements achieved by each of these components.
Ours CoScene Splits on SemanticKITTI Avg.
PtErase MixPatch InsFill 1% 10% 20% 50%
45.4 57.1 59.2 60.0 55.4
51.7 59.8 61.8 60.4 58.4
51.3 61.4 60.7 60.5 58.5
52.3 61.7 60.1 61.1 58.8
54.6 61.0 62.9 62.8 60.3
52.0 61.7 60.8 62.6 59.3
50.5 62.5 63.8 63.6 60.1
54.5 63.3 63.7 64.3 61.5
Table 2: Ablation study of different components of our designed CoScene. PtErase refers to the point erause strategy, while MixPatch and InsFill refers to the patch mix and instance filling data augmentation, respectively. All mIoU results are reported in percentage (%).
Figure 5: Performance comparison on SemanticKITTI val split under different semi-supervised settings. (a) Illustration of performance differences whether using consistency assurance strategy on the LaserMix (Kong et al. 2023b). (b) Comparison between dual-scene level and multi-scene level for the patch-based data augmentation in our CoScene network. (c) Results of mIoU against the number of patch split $n$ under the 10% SemanticKITTI split.
Effectiveness of point erasure strategy. The row 1 in Table 2 denotes the performance of MeanTeacher (Tarvainen and Valpola 2017) as the baseline, revealing a marginal enhancement compared to the supervision-only Cylinder3D model (Zhu et al. 2021). When PtErase is applied to the baseline, it is effective and can boost the performance by ensuring consistency during forward and backward propagation, thereby compensating for the shortcomings of the pseudo-label mechanism. The consistency assurance is implemented through a plug-and-play point erasure strategy, which can be easily integrated into methods based on the pseudo-label mechanism. To further validate the significance of ensuring consistency during forward and backward propagation, we embedded the PtErase strategy into LaserMix. As illustrated in Figure 5 (a), performance improves across all ratios when using the PtErase strategy. This further emphasizes the importance of intra-scene consistency and demonstrates the generalization ability of our method.
Effectiveness of patch-based data augmentation. The patch-based data augmentation consists of MixPatch (patch-based mix) and InsFill (instance filling). Rows 3 and 4 of Table 2 demonstrate that using MixPatch or InsFill alone results in an average mIoU improvement of 3.1% or 3.4%, respectively, compared to the baseline. This suggests that annotation information transfer across scenes benefits SSL 3D segmentation. As shown in rows 5 and 6 of the Table 2, combining MixPatch or InsFill with PtErase significantly enhances performance, which indicates the effectiveness of the patch-based data augmentation. Furthermore, the collaboration of all three components, leads to substantial performance gains, as validated by the last row of Table 2. Additionally, our patch pool and instance pool within the patch-based data augmentation facilitate effective information fusion across multiple scenes. This approach inherently enables the transfer of inter-scene diverse correlation information. To verify the effectiveness of the interaction at the multi-scene level, we modify the pool implementation to enable only interaction between two scenes. The comparative results, as shown in Figure 5 (b), indicate superior performance at multi-scene level compared to dual-scene level.
Different mixing strategies. Figure 5 (c) compares the patch-based data augmentation employed in our CoScene with other mixing approaches (Nekrasov et al. 2021; Xiao et al. 2022; Yun et al. 2019; Zhang et al. 2017). MixUp demonstrates a slight improvement over the baseline, likely due to its limited ability to transform information. Other methods have a considerable improvement over the baseline by enriching point cloud distributions. In contrast, our CoScene achieves the greatest gain by effectively exploring rich inter-scene correlation, highlighting its effectiveness in facilitating information transfer between labeled and unlabeled scenes. This is accomplished through its innovative design, which includes a patch pool and an instance pool spanning multiple scenes.
Patch split number. As the patch split number $n$ plays a critical role in our framework, we further analyze the impact of varying $n$ in patch-based data augmentation, with the results presented in Figure 6. When $n$ is small, it can yield a certain level of enhancement, but it may fail to generate data sufficiently divergent from the original, impeding the effective transfer of information between labeled and unlabeled scenes. Conversely, when $n$ is too large, the generated scenes may deviate excessively from the original data, resulting in the loss of essential prior scene information. Empirically, we identify that setting $n$ to 18 yields optimal results. Therefore, throughout this paper, we consistently use $n$ set to 18, unless explicitly stated otherwise.
Figure 6: Comparison of different mixing-based techniques.
Conclusion
In this paper, we explore the scene coherence for semi-supervised point cloud segmentation, building upon the intra-scene consistency and inter-scene correlation. In particular, we achieve intra-scene consistency assurance through a point erasure strategy that ensures consistency during the forward and backward propagation, rather than simply disregarding points with pseudo-labels as in previous methods. Additionally, we explore inter-scene correlation and propose a patch-based data augmentation to enhance information transfer between labeled scenes and unlabeled scenes. Benefiting from these two components, our CoScene effectively transfers diverse semantic information and produces more accurate segmentation results. Extensive experimental results demonstrate the superiority of our CoScene over state-of-the-art methods.

### v2

@@
-Exploring Scene Coherence for Semi-Supervised 3D Semantic Segmentation
+Exploring Scene Affinity for Semi-Supervised LiDAR Semantic Segmentation
 Abstract
-Semi-supervised semantic segmentation, which efficiently addresses the limitation of acquiring dense annotations, is essential for 3D scene understanding.
-Most methods leverage the teacher model to generate pseudo labels, and then guide the learning of the student model on unlabeled scenes.
-However, they focus only on points with pseudo labels while directly overlooking points without pseudo labels, namely intra-scene inconsistency, leading to semantic ambiguity.
-Moreover, inter-scene correlation between labeled and unlabeled scenes contribute to transferring rich annotation information, yet this has not been explored for the semi-supervised tasks.
-To address these two problems, we propose to explore scene coherence for semi-supervised 3D semantic segmentation, dubbed CoScene.
-Inspired by the unstructured and unordered nature of the point clouds, our CoScene adopts the straightforward point erasure strategy to ensure the intra-scene consistency.
-Moreover, patch-based data augmentation is proposed to enhance the inter-scene information transfer between labeled and unlabeled scenes at both scene and instance levels.
-Extensive experimental results on SemanticKITTI and nuScenes show that our approach outperforms existing methods.
-Introduction
-LiDAR (Light Detection and Ranging) point cloud semantic segmentation, aiming to recognize the class label for each point of an input point cloud scene, is a core task in autonomous driving.
-Due to the advancement of deep neural networks, tremendous progress has been achieved (Hu et al.
-2020; Zhu et al.
-2021; Su et al.
-2023).
-However, the excellent performance of most studies depends heavily on large-scale datasets with high-quality point-level annotations.
-Such dense and delicate point-level labeling is labor-intensive and time-consuming, hindering the deployment of point cloud semantic segmentation models in real-world applications.
-Therefore, there is a compelling need to reduce the dependence on huge exhaustively labeled datasets.
-Figure 1: (a) Generated pseudo-label scenes from unlabeled scenes containing points with and without pseudo labels simultaneously (green and gray).
-(b) Regardless of whether point-based or voxel-based segmentation networks are employed, resulting in inconsistencies between forward and backward propagation.
-(c) Annotation information transfer between labeled and unlabeled scenes.
-Semi-supervised learning (SSL), a significant machine learning paradigm for alleviating data annotation demands, seeks to utilize a small number of labeled samples alongside a vast collection of unlabeled samples to optimize deep models.
-Recently, several semi-supervised 3D semantic segmentation methods (Kong et al.
-2023b; Jiang et al.
-2021; Li, Shum, and Breckon 2023) have been proposed, with a core focus on effectively transferring diverse semantic information.
-These methods typically first obtain pseudo-labeled scenes from unlabeled data using the teacher network, and then feed the entire pseudo-labeled scenes into the student network for training.
-Although these methods have reported promising results, two main issues still need to be addressed.
-As shown in Figure 1 (a), the pseudo-labeled scene consists of points with and without pseudo-labels simultaneously.
-During training, all points are involved in the model’s forward propagation, while only points with pseudo-labels are used for backward propagation, and other points are overlooked.
-Inconsistency between forward and backward propagation within the scene leads to semantic ambiguity, meaning the segmentation model is unable to determine the semantic labels for points without pseudo-labels.
-Consequently, this limits the effectiveness of semantic information transfer.
-Both point-based (Hu et al.
-2020) and voxel-based (Choy, Gwak, and Savarese 2019; Zhu et al.
-2021) methods suffer from this issue, as illustrated in Figure 1 (b).
-Diverse semantic information is crucial for the success of semi-supervised learning of the segmentation models (Kong et al.
-2023b; Xiao et al.
-2022).
-Thus, many studies focus on developing data augmentation methods (Sheshappanavar, Singh, and Kambhamettu 2021; Zhang et al.
-2022) to expand the data distribution.
-However, these methods typically model the correlation between two scenes, while neglecting the correlation among multiple scenes.
-As shown in Figure 1 (c), the labeled and unlabeled scenes may exhibit similarities in factors such as spatial structures.
-Exploring the correlations between scenes can be beneficial for annotation information transfer, thereby enhancing learning on unlabeled scenes.
-To address the above two issues, we explore Scene Coherence and propose CoScene for semi-supervised 3D semantic segmentation.
-Specifically, based on the unstructured and unordered nature of point clouds, we design a point erasure strategy that removes the points without pseudo labels during forward propagation, achieving full annotation of the training scene and ensuring intra-scene consistency.
-This strategy complements the insufficiency of the basic pseudo-label mechanism, which can serve as a plug-and-play module for semi-supervised 3D segmentation frameworks based on pseudo labels.
-Additionally, we design a patch-based data augmentation to model the correlation between multiple scenes at both the scene and instance levels.
-The key idea is to construct two patch pools to store scene-level and instance-level information, and then combine scene-level and instance-level patches from multiple scenes to generate diverse training samples, facilitating the transfer of annotation information to unlabeled scenes.
-The proposed CoScene has been extensively evaluated through experiments and has demonstrated superiority across various benchmarks for semi-supervised 3D semantic segmentation.
-Under the setting of 5% labeled scenes, our method outperforms the state-of-the-art method by 4% in terms of mIoU on the SemanticKITTI (Behley et al.
-2019) dataset.
+This paper explores scene affinity (AIScene), namely intra-scene consistency and inter-scene correlation, for semi-supervised LiDAR semantic segmentation in driving scenes.
+Adopting teacher-student training, AIScene employs a teacher network to generate pseudo-labeled scenes from unlabeled data, which then supervise the student network’s learning.
+Unlike most methods that include all points in pseudo-labeled scenes for forward propagation but only pseudo-labeled points for backpropagation, AIScene removes points without pseudo-labels, ensuring consistency in both forward and backward propagation within the scene.
+This simple point erasure strategy effectively prevents unsupervised, semantically ambiguous points (excluded in backpropagation) from affecting the learning of pseudo-labeled points.
+Moreover, AIScene incorporates patch-based data augmentation, mixing multiple scenes at both scene and instance levels.
+Compared to existing augmentation techniques that typically perform scene-level mixing between two scenes, our method enhances the semantic diversity of labeled (or pseudo-labeled) scenes, thereby improving the semi-supervised performance of segmentation models.
+Experiments show that AIScene outperforms previous methods on two popular benchmarks across four settings, achieving notable improvements of 1.9% and 5.3% in the most challenging 1% labeled data.
+1 Introduction
+Semantic segmentation of LiDAR point clouds is essential for autonomous driving, allowing vehicles to perceive their 3D surroundings.
+Most state-of-the-art models for driving-scene LiDAR segmentation are dependent on high-quality, point-annotated datasets [12, 55, 35].
+With numerous objects in driving scenes, densely annotating points is extremely labor-intensive.
+Consequently, semi-supervised LiDAR semantic segmentation has attracted growing attention, where models are trained on a limited amount of labeled data supplemented by a large set of unlabeled data.
+Figure 1: Top: Pseudo-labeled scenes generated from unlabeled scenes typically contain points with pseudo labels (colored) and without them (gray).
+Previous methods include all points in forward propagation but only pseudo-labeled points in backward propagation, leading to intra-scene inconsistency.
+This inconsistency may allow unsupervised, semantically ambiguous points to affect the learning of pseudo-labeled points (see Fig.
+7).
+Bottom: Existing data augmentation techniques suitable for semi-supervised segmentation usually mix two scenes through scene-level operations like concatenation or swapping.
+Their resulting scenes may be constrained in terms of semantic diversity, potentially lacking objects like bicyclists and buses, which are important in driving scenes.
+The common solution for learning from unlabeled data is to apply teacher-student training, where a teacher network generates pseudo-labeled scenes from unlabeled data, which are then used to supervise the learning of a student network.
+The student network’s objective is to minimize segmentation loss while the teacher network’s weights are updated via exponential moving average [38].
+To further improve model learning with unlabeled data, existing studies implement designs such as utilizing language-vision models [22], adopting density-guided contrastive learning [21], or trying feature fusion between LiDAR and image data [5].
+Despite the progress, two main issues remain to be addressed:
+(1) Given the teacher network’s confident prediction, pseudo-labeled scenes typically contain points with and without pseudo-labels.
+During training, most studies include all points in pseudo-labeled scenes in the student network’s forward propagation, but only pseudo-labeled points are involved in backward propagation.
+Points excluded from backpropagation are unsupervised and, therefore, semantically ambiguous.
+However, these points still contribute to the forward feature computation for pseudo-labeled points, which may interfere with the model’s ability to learn from them effectively.
+In short, the inconsistency between forward and backward propagation within the scene causes ambiguous points to obstruct the model’s learning of pseudo-labeled points.
+Both point-based [12] and voxel-based [7, 55] methods suffer from this issue in feature computation, as illustrated at the top of Fig.
+1.
+(2) Data augmentation is an effective practice in semi-supervised LiDAR semantic segmentation [15, 13].
+The basic idea is to mix raw labeled scenes and pseudo-labeled scenes to enrich scene semantics.
+Although impressive advancements have been made [32, 50], they usually select two scenes (one labeled scene and one pseudo-label scene) and then perform scene-level mix through operations like concatenation [26] and swapping [15], which limits the semantic diversity of mixed scenes.
+As shown at the bottom of Fig.
+1, multiple scenes, whether labeled or pseudo-labeled, often share similar spatial layouts.
+Moreover, in driving scenes, objects such as cars and pedestrians are the primary focus of an autonomous vehicle’s attention.
+This naturally raises the question: can we improve the semantic diversity of scenes by mixing multiple scenes at the instance level?
+To address inconsistency and implement the multi-scene mixing idea, we explore Scene AffInity for semi-supervised LiDAR semantic segmentation, named AIScene.
+The term scene affinity refers to intra-scene consistency and inter-scene correlation.
+Specifically, AIScene, built on the teacher-student training pipeline, employs a simple yet helpful point erasure strategy, removing points without pseudo-labels in forward propagation.
+This ensures intra-scene consistency in both forward and backward propagation, preventing semantically ambiguous points from disrupting learning from pseudo-labeled points.
+Building on this, AIScene incorporates patch-based data augmentation that mixes multiple scenes at both scene and instance levels.
+By maintaining two pools that store scene and instance patches from various scenes, it selects patches for mixing based on inter-scene correlation.
+This novel augmentation method not only enhances the semantic diversity of raw labeled scenes but also complements pseudo-labeled scenes where some points have been erased.
+AIScene has been extensively evaluated through experiments and demonstrated superiority on two popular benchmarks for semi-supervised LiDAR semantic segmentation.
+It exceeds state-of-the-art methods by 1.3% on the SemanticKITTI [2] and 2.4% on the nuScenes [10] in terms of average performance across four settings (1%, 10%, 20%, and 50% labeled data).
+Notably, in the most challenging setting, with only 1% labeled data, AIScene achieves remarkable improvements of 1.9% and 5.3% on the two datasets.
+Additionally, our ablation studies confirm AIScene’s potential to reduce the requirements for labeled data across different LiDAR semantic segmentation networks (see Tab.
+1 and  Tab.
+2), with each component of AIScene able to function independently to facilitate semi-supervised LiDAR semantic segmentation (see Fig.
+6).
 To summarize, our contributions are as follows:
-  • We propose a plug-and-play point erasure strategy complementing with pseudo-label mechanism in semi-supervised point cloud semantic segmentation, ensuring intra-scene consistency.
-  • We propose a patch-based data augmentation to model inter-scene correlation information at scene and instance levels simultaneously, facilitating information transfer between labeled scenes and unlabeled scene.
-  • Extensive experiments demonstrate that our semi-supervised 3D segmentation methods achieves state-of-the-art performance, and comprehensive ablations studies are conducted to demonstrate its effectiveness.
-Figure 2: Overview of the proposed CoScene training pipeline.
+  • We propose a point erasure strategy to ensure intra-scene consistency and facilitate network learning from pseudo-labeled scenes.
+This strategy can serve as a plug-and-play tool for semi-supervised LiDAR semantic segmentation using the pseudo-label mechanism.
+  • We design a patch-based data augmentation to mix multiple scenes at both scene and instance levels, guided by inter-scene correlation.
+This method improves the semantics diversity of labeled scenes and complements pseudo-labeled scenes.
+  • Extensive experiments demonstrate that AIScene achieves state-of-the-art results on popular benchmarks for semi-supervised LiDAR semantic segmentation.
+Ablation studies show that the novel erasure strategy and augmentation method can function independently.
+Figure 2: Overview of the proposed AIScene training pipeline.
 The teacher network generates the pseudo-labels $y_{i}^{u}$ from the unlabeled scene $x_{i}^{u}$, followed by an erasure operation to filter out low-confidence points, ensuring training consistency.
-For the labeled scene $x_{i}^{l}$ and the generated erased scene $\hat{x}_{i}^{u}$, we employ a patch-based data augmentation to obtain augmented scenes $x_{i}^{lpf}$ and $\hat{x}_{i}^{upf}$, separately.
-Subsequently, both scenes are fed into the student network for computing the loss of $\mathcal{L}_{l}$ and $\mathcal{L}_{u}$.
-Related Work
-LiDAR Semantic Segmentation
-LiDAR semantic segmentation aims to recognize the class label for each point of an input point cloud scene.
-Through the development of various deep neural networks, numerous approaches for LiDAR semantic segmentation have been explored over the past few years, which can be broadly grouped into three categories, i.e., voxel-based (Zhu et al.
-2021), 2D projection-based (Kong et al.
-2023a; Zhao, Bai, and Huang 2021), and point-based methods (Hu et al.
-2020; Qi et al.
-2017).
-Prior fully supervised works have made significant progress and achieved appealing performance, while the results deeply depend on large-scale precisely annotated datasets (Behley et al.
-2019; Fong et al.
-2022).
-To reduce annotations of 3D data, some methods with less supervision have been explored recently, e.g., weakly-supervised (Pan et al.
-2024; Liu et al.
-2022; Unal, Dai, and Van Gool 2022; Zhang et al.
-2021; Chen et al.
-2023; Tang et al.
-2024) and semi-supervised learning (Li, Shum, and Breckon 2023; Reichardt, Ebert, and Wasenmüller 2023; Zou et al.
-2018; Chen et al.
-2021).
-LaserMix (Kong et al.
-2023b), as the pioneer in semi-supervised LiDAR segmentation leverages the simple laser beam mix strategy to get solid results.
-Follow-up work fuses LiDAR-Camera data to mine deeper information (Chen et al.
-2024).
-For semi-supervised approaches, the overall architecture often builds upon the mean-teacher paradigm (Tarvainen and Valpola 2017) and leverages the concept of pseudo label (Arazo et al.
-2020).
-However, many points without pseudo labels contribute equally to pseudo-label points in the forward propagation, but are directly ignored when performing backward propagation, which brings intra-scene inconsistency issues.
-Hence, our CoScene leverages the point erasure strategy to mitigate this inconsistency problem, creating a complement to the pseudo-label mechanism to enable effective information transfer.
-3D Point Cloud Data Augmentation
-Point cloud data augmentation has been explored across various 3D vision tasks (Zhu, Fan, and Weng 2023; Shin et al.
-2023; Wang et al.
-2024; Fang et al.
-2021).
-Enlightened by the success of Mixup (Zhang et al.
-2017) in image classification, several works explore the techniques for mixing selected point cloud samples.
-PointMixup (Chen et al.
-2020) leverages interpolation to generate mixed samples for training.
-To preserve structural information, RSMix (Lee et al.
-2021) provides a neighboring function, and PointCutMix (Zhang et al.
-2022) establishes the correspondence relationship for sample fusion.
-Notably, these works aim at local object-level augmentation.
-As a result, several works also study the mixing strategies for point cloud scene augmentation.
-Mix3D (Nekrasov et al.
-2021) concatenates points of two scenes to create a new training sample.
-PloarMix (Xiao et al.
-2022) swaps points within the fan-shaped area between two scenes.
-CoSMix (Saltori et al.
-2023) aims to mitigate the domain shift by mixing semantic point groups.
-These works only support similarity modeling between two scenes, but fail to construct correlation among multiple scenes.
-By contrast, our approach has the capability to model inter-scene information from multiple scenes to create more diversified data, while also extracting instance-level semantic information for individual classes to further enhance scenes.
-Therefore, our approach can effectively guide the transfer of annotation information from labeled scenes to unlabeled scenes, facilitating semi-supervised learning.
-Method
-In this section, we first state the definition of semi-supervised 3D semantic segmentation and introduce the mainstream teacher-student framework.
-Then, we present the intra-scene consistency and inter-scene correlation of CoScene, along with its core component, i.e., the point erasure strategy and patch-based data augmentation.
-Finally, we outline the overall training objective.
-Preliminary
+For the labeled scene $x_{i}^{l}$ and the generated erased scene $\hat{x}_{i}^{u}$, we employ patch-based data augmentation to obtain augmented scenes $x_{i}^{lpf}$ and $\hat{x}_{i}^{upf}$, respectively.
+Subsequently, both scenes are fed into the student network to compute the loss of $\mathcal{L}_{l}$ and $\mathcal{L}_{u}$.
+2 Related Work
+Fully-supervised LiDAR Semantic Segmentation: Numerous approaches for LiDAR semantic segmentation have been explored, which can be broadly grouped into three categories: voxel-based, 2D projection-based, and point-based methods.
+Voxel-based approaches [6, 7, 36, 55, 19] involve voxelizing sparse point clouds to generate regular voxel grids, allowing the application of standard 3D convolution neural networks for embedding learning.
+Conversely, 2D projection-based methods [14, 25, 43, 46, 51, 53] project input point clouds into a 2D dense space, use 2D convolutional filters [30] for processing, and then map predictions back to the original 3D space.
+Finally, point-based methods [12, 28, 18, 44] operate directly on raw point clouds, enabling the extraction of point-wise semantic features.
+Semi-supervised LiDAR Semantic Segmentation: To reduce the need for 3D dense annotations, recent research has explored methods that require less supervision [17, 24, 39, 52, 3, 37].
+Semi-supervised learning [23, 29, 56, 4, 21, 16] aims to utilize a large number of unlabeled data to alleviate the annotation burden.
+LaserMix [15], a pioneer work in semi-supervised LiDAR segmentation, leverages a simple laser beam mix strategy to achieve solid results.
+Subsequent research fuses LiDAR and camera data to extract deeper semantic information [5, 40].
+DDSemi [21], focusing on point cloud density, builds on contrastive learning to enhance feature discriminability.
+The recently proposed LASS3D [22] leverages large language-vision models to support semi-supervised segmentation.
+Existing semi-supervised methods often build upon the mean-teacher paradigm [38] and utilize pseudo labels [1].
+However, many points without pseudo-labels contribute equally to pseudo-labeled points in forward propagation but are ignored during backward propagation, leading to intra-scene inconsistency.
+To address this, our AIScene leverages a point erasure strategy to mitigate this inconsistency, complementing the pseudo-label mechanism.
+3D Point Cloud Data Augmentation: Point cloud data augmentation has been explored across various 3D vision tasks [54, 33, 42, 9].
+Inspired by the success of Mixup [49] in image classification, several methods have investigated techniques for mixing selected point cloud samples.
+To preserve structural information, RSMix [20] introduces a neighboring function, and PointCutMix [50] establishes correspondences for sample fusion.
+Notably, these methods focus on local object-level augmentation.
+In response, several works have explored mixing strategies for point cloud scene augmentation.
+PloarMix [45] swaps points within the fan-shaped areas between two scenes, while CoSMix [31] aims to mitigate domain shift by mixing semantic point groups.
+Unlike these methods, our method is capable of modeling inter-scene information across multiple scenes to create more diversified data, while also extracting instance-level semantic information for individual classes to further enrich scenes.
+3 Method
+3.1 Preliminary
 Problem Formulation.
@@
 In detail, we have a small labeled training set $D_{l}=\left\{x_{i}^{l},y_{i}^{l}\right\}_{i=1}^{N_{l}}$ and a large unlabeled training set $D_{u}=\left\{x_{i}^{u}\right\}_{i=1}^{N_{u}}$, where $x_{i}\in\mathbb{R}^{m\times{(3+r)}}$ is a point cloud scene of $m$ points with three-dimensional Cartesian coordinates relative to the sensor and additional $r$-dimensional feature (e.g., intensity, color, or reflectivity), $y_{i}$ represents its corresponding point-level segmentation labels, $N_{l}$ and $N_{u}$ are the numbers of labeled and unlabeled point cloud scenes, respectively, and $N_{u}\gg N_{l}$ in ideal cases.
-The goal of SSL-based point cloud segmentation is to learn a segmentation model using the limited labeled data and large-scale unlabeled data.
+The goal of SSL-based point cloud segmentation is to learn a segmentation model using limited labeled data and large-scale unlabeled data.
 Teacher-Student Paradigm.
-Following current mainstream research (Li, Shum, and Breckon 2023), our learning framework also builds upon the teacher-student paradigm, which includes two 3D segmentation networks with the same architecture.
-Here, we adopt Cylinder3D (Zhu et al.
-2021) as the segmentation backbone $\Phi(\cdot)$ to enable a fair comparison with other methods.
-In detail, the teacher model generates pseudo-labels for unlabeled scenes and undergoes a gradual update through the exponential moving average (EMA) (Tarvainen and Valpola 2017) of the student weights:
-$\theta_{\mathbb{t}}^{k+1}=\alpha\cdot\theta_{\mathbb{t}}^{k}+(1-\alpha)\cdot\theta_{\mathbb{s}}^{k},$ (1)
-where $\alpha$ is the EMA decay rate, which is set to 0.99 by default, $\theta_{\mathbb{t}}$ and $\theta_{\mathbb{s}}$ represent the weight parameters of the teacher and student models, respectively, and $k$ denotes the training step.
-Point Erasure Strategy
-Based on the teacher-student mutual learning framework (Sohn et al.
-2020), the overall architecture of CoScene is depicted in Figure 2.
-Specifically, CoScene consists of two primary branches: one dedicated to labeled scenes and the other to unlabeled scenes.
+Following current mainstream research [15, 21], our framework also builds upon the teacher-student paradigm, which includes two segmentation networks with the same architecture.
+Our AIScene can be instantiated with any 3D segmentation network [55, 7, 19], denoted as $\Phi(\cdot)$.
+In detail, the teacher model generates pseudo-labels for unlabeled scenes and undergoes a gradual update through the exponential moving average (EMA) [38] of the student weights:
+$\theta_{\mathbbm{t}}^{k+1}=\alpha\cdot\theta_{\mathbbm{t}}^{k}+(1-\alpha)\cdot\theta_{\mathbbm{s}}^{k},$ (1)
+where $\alpha$ is the EMA decay rate, which is set to 0.99 by default, $\theta_{\mathbbm{t}}$ and $\theta_{\mathbbm{s}}$ represent the weight parameters of the teacher and student models, respectively, and $k$ denotes the training step.
+3.2 Point Erasure Strategy
+Based on the teacher-student mutual learning framework [34], the overall architecture of AIScene is depicted in Fig.
+2.
+Specifically, AIScene consists of two primary branches: one dedicated to labeled scenes and the other to unlabeled scenes.
 How to effectively leverage a large number of unlabeled scenes is crucial for semi-supervised learning.
-For the branch of unlabeled scenes, we use a teacher network to generate pseudo-labeled scenes $(x_{i}^{u},y_{i}^{u})$ based on a predefined confidence threshold $\tau_{s}$ (Kong et al.
-2023b).
-As mentioned previously, when using the pseudo-labeled scene for training, points without pseudo labels are involved in the forward propagation but are ignored in the backward propagation due to the lack of corresponding supervision.
-This leads to intra-scene inconsistency, which limits the effectiveness of semantic information transfer.
+Hence, for the branch of unlabeled scenes, we use a teacher network to generate pseudo-labeled scenes $(x_{i}^{u},y_{i}^{u})$ based on a predefined confidence threshold $\tau_{s}$ [15].
+As mentioned previously, when using the pseudo-labeled scene for training, points without pseudo-labels are involved in the forward propagation but are ignored in the backward propagation due to the lack of corresponding supervision.
+This leads to intra-scene inconsistency, which causes semantic ambiguity of points.
 To address this issue, we propose erasing points without pseudo labels, resulting in an erased scene $(\hat{x}_{i}^{u},\hat{y}_{i}^{u})$ that contains fully annotated information, similar to fully supervised scenes.
@@
 where $\hat{x}_{i}^{u}$ is the erased scene with corresponding class labels $\hat{y}_{i}^{u}$, $\Phi_{s}(\cdot)$ represents the segmentation network that predicts the confidence score for each point in the scene $x_{i}^{u}$, and $\tau_{s}$ is a predefined confidence threshold, equivalent to the pseudo-label mining threshold, requiring no additional adjustment.
-Since the proposed point erasure strategy is independent of the segmentation network architecture, it can be utilized as a plug-and-play tool for semi-supervised 3D semantic segmentation methods based on the pseudo-label mechanism (see Figure 5 (a)).
+Since the proposed point erasure strategy is independent of the segmentation network architecture, it can be utilized as a plug-and-play tool for semi-supervised 3D semantic segmentation methods based on the pseudo-label mechanism, as depicted in Fig.
+6 (a).
 Figure 3: Illustration of the pool initialization process.
-For each labeled point cloud scene, splitting it into multiple patches from the BEV perspective and assigning different instance categories to each patch enable the extraction of instances belonging to different classes.
+For each labeled point cloud scene, splitting it into multiple patches from the BEV perspective and assigning different instance categories to each patch enables the extraction of instances belonging to different classes.
 Subsequently, these patches and instances are stored individually in the patch pool and instance pool.
-Patch-based Data Augmentation
-Based on the labeled and erased scenes, we propose a patch-based data augmentation to enhance the transfer of annotation information, guided by inter-scene correlation.
+3.3 Patch-based Data Augmentation
+Based on the labeled and erased scenes, we propose a patch-based data augmentation to enhance the semantic diversity of annotation information, guided by inter-scene correlation.
 The key idea is to construct patch pools to store scene-level and instance-level information, and then combine patches from multiple scenes to generate diverse training samples.
 Pool Initialization.
-We first use the labeled training set $D_{l}$ to create two patch pools, which store local-scene and single-instance information, respectively, as shown in Figure 3.
+We first use the labeled training set $D_{l}$ to create two patch pools, which store scene-level and instance-level information, respectively, as shown in Fig.
+3.
 Given the $i$-th point cloud scene $x_{i}^{l}$, we patchify $x_{i}^{l}$ into a sequence of $T$ regular, non-overlapping point patches $\{p_{ij}^{l}\}_{j=1}^{T}$, each of size $(n\times n)$, in the bird’s eye view.
@@
 $\left\{p_{ij}^{l}\right\}_{j=1}^{T}=\text{patchify}(x_{i}^{l}),$ (3)
-where $p_{ij}^{l}$ denotes the $j$-th point patch from the sequence. From all labeled scenes, we then construct the patch pool $\mathcal{B}_{\text{Pat}}^{l}$, which can be defined as:
+where $p_{ij}^{l}$ denotes the $j$-th point patch from the sequence.
+For all labeled scenes, by perform the operation in Eq.
+3, the patch pool $\mathcal{B}_{\text{Pat}}^{l}$ is constructed, which can be defined as:
 $\mathcal{B}_{\text{Pat}}^{l}=\left\{\left\{p_{ij}^{l}\right\}_{j=1}^{T}\right\}_{i=1}^{N_{l}}.$ (4)
-Based on the patch pool, we can obtain the instance pool $\mathcal{B}_{\text{Ins}}^{l}$ by assigning different class labels to each patch, which contains instance information for different categories (e.g., car and person).
+Based on the patch pool, we can obtain the instance pool $\mathcal{B}_{\text{Ins}}^{l}$ by assigning different class labels to each patch, as depicted in Fig.
+3, which contains instance information for different categories (e.g., car and person).
 Following the same steps used for labeled scenes, we further create a pseudo patch pool $\mathcal{B}_{\text{Pat}}^{u}$ and a pseudo instance pool $\mathcal{B}_{\text{Ins}}^{u}$ from the erased scene $\hat{x}_{i}^{u}$.
@@
 This prevents low-quality pseudo labels from persisting in pools and misleading model training.
-Additionally, if the number of points within a patch or instance falls below a certain threshold $\tau_{\text{min}}$, it will not be saved into the pool, ensuring effective inter-scene information transfer.
+Additionally, if the number of points within a patch or instance falls below a certain threshold $\tau_{\text{min}}$, it will not be saved into the pool, ensuring effective semantic diversity.
 Figure 4: (a) Illustration of the MixPatch process.
@@
 MixPatch.
-To model the inter-scene correlation and then guide the transfer of annotation information between labeled scenes and unlabeled scenes, we combine multiple scenes using patch pools, as shown in Figure 4.
-Figure 4 (a) illustrates the patch mixing (MixPatch), while Figure 4 (b) presents the instance filling (InsFill), which is introduced in the next paragraph.
+To model the inter-scene correlation and then guide the transfer of annotation information between labeled scenes and unlabeled scenes, we combine multiple scenes using patch pools.
+Fig.
+4 (a) illustrates the patch mixing (MixPatch) strategy, while Fig.
+4 (b) presents the instance filling (InsFill) strategy, which is introduced in the next paragraph.
 During a training iteration, the mini-batch is comprised of labeled and unlabeled scenes.
-After the unlabeled scene $x_{i}^{u}$ is transformed into the erased scene $\hat{x}_{i}^{u}$ using the point erasure strategy, we integrate information from the patch pool into $\hat{x}_{i}^{u}$.
-The patch mixing process can be described as follows:
+After the unlabeled scene $x_{i}^{u}$ is transformed into the erased scene $\hat{x}_{i}^{u}$ using the point erasure strategy, we integrate information from the patch pool into $\hat{x}_{i}^{u}$ by MixPatch.
+The MixPatch process can be described as follows:
 $\hat{x}_{i}^{up}=(\hat{x}_{i}^{u}\odot\mathcal{M})\oplus(\mathcal{B}_{\text{Pat}}^{l}\odot(1-\mathcal{M})),$ (5)
@@
 where $\mathcal{M}$ is the same as defined in Eq.
-(5), thereby ensuring the full utilization of scene information from the unlabeled scenes.
-The binary mask $\mathcal{M}$ and its complement $1-\mathcal{M}$ are used to combine inter-scene point cloud patches with the erased scene, facilitating the transfer of annotation information between labeled and unlabeled scenes.
-To alleviate bias in the patch mixing process, we apply a uniform sampling strategy on the patches of each scene, ensuring that each patch has an equal opportunity to be selected in the mixed scene.
+5, thereby ensuring the full utilization of scene information from the unlabeled scenes.
+The binary mask $\mathcal{M}$ and its complement $1-\mathcal{M}$ are used to merge inter-scene point cloud patches with the erased scene, facilitating the generation of diverse semantic information for labeled and unlabeled scenes.
+To alleviate bias in the patch mixing process, we apply a uniform sampling strategy to the patches in each scene, ensuring that each patch has an equal opportunity of being selected in the mixed scene.
 InsFill.
@@
 Given a patch with location index $j$ in the scene, we randomly select an instance from the instance pool $\mathcal{B}_{\text{Ins}}^{l}$ that was extracted from the $j$-th patch.
-Patches with identical indices typically correspond to similar LiDAR spatial positions, ensuring consistency between the selected instance and its fill positions in the scene.
-As depicted in Figure 4 (b), if the selected instance overlaps with existing instances in the scene or lacks sufficient contextual information, the filling operation is bypassed.
+Patches with identical indices typically correspond to similar LiDAR spatial layouts, ensuring consistency between the selected instance and its fill positions in the scene.
+As depicted in Fig.
+4 (b), if the selected instance overlaps with existing instances in the scene or lacks sufficient contextual information, the filling operation is bypassed.
 This judgment process can be easily implemented using the maximum 3D bounding boxes.
-This data augmentation provides single-instance semantic information for the point cloud scene, enabling the augmented scene $\hat{x}_{i}^{upf}$ and corresponding labels $\hat{y}_{i}^{upf}$ to effectively support semi-supervised training.
+This data augmentation provides instance-level semantic information for the point cloud scene, enabling the augmented scene $\hat{x}_{i}^{upf}$ and corresponding labels $\hat{y}_{i}^{upf}$ to effectively support semi-supervised training.
 For the labeled branch, we also apply the same filling strategy as described, selecting instances from the pseudo instance pool $\mathcal{B}_{\text{Ins}}^{u}$ to fill the current scene.
-This process yields the scene $x_{i}^{lpf}$ along with its corresponding label $y_{i}^{lpf}$.
-Model Training
-Following the standard teacher-student mutual learning paradigm (Kong et al.
-2023b; Sohn et al.
-2020), we construct a mini-batch of data with an equal composition of labeled and unlabeled scenes at each iteration.
+This InsFill process yields the scene $x_{i}^{lpf}$ along with its corresponding label $y_{i}^{lpf}$.
+Method Venue SemanticKITTI [2] Avg. nuScenes [10] Avg.
+1% 10% 20% 50% 1% 10% 20% 50%
+Meanteacher [38] NeurIPS’2017 45.4 57.1 59.2 60.0 55.4 51.6 66.0 67.1 71.7 64.1
+CBST [56] ECCV’2018 48.8 58.3 59.4 59.7 56.6 53.0 66.5 69.6 71.6 65.2
+CPS [4] CVPR’2021 46.7 58.7 59.6 60.5 56.4 52.9 66.3 70.0 72.5 65.4
+LaserMix (Range View) [15] CVPR’2023 43.4 58.8 59.4 61.4 55.8 49.5 68.2 70.6 73.0 65.3
+LaserMix (Voxel) [15] CVPR’2023 50.6 60.0 61.9 62.3 58.7 55.3 69.9 71.8 73.2 67.6
+GPC [13] ICCV’2021 54.1 62.0 62.5 62.8 60.4 - - - - -
+IGNet [40] WACV’2024 49.0 61.3 63.1 64.8 59.6 - - - - -
+LiM3D [23] CVPR’2023 58.4 62.2 63.1 63.6 61.8 - - - - -
+LASS3D [22] ECCV’2024 58.5 63.0 64.1 64.5 62.5 - - - - -
+DDSemi [21] CVPR’2024 59.3 65.1 66.3 67.0 64.4 58.1 70.2 74.0 76.5 69.7
+Ours (AIScene) - 61.2 66.3 67.4 67.9 65.7 63.4 73.7 75.0 76.3 72.1
+Improvement $\uparrow$ - +1.9 +1.2 +1.1 +0.9 +1.3 +5.3 +3.5 +1.0 -0.2 +2.4
+Table 1: Experimental results on the SemanticKITTI and nuScenes dataset compared with different SSL methods.
+The results are reported with mIoU scores (%).
+The blod and underline scores respectively denote the best and second results.
+3.4 Model Training
+Following the standard teacher-student mutual learning paradigm [38, 15, 34], we construct a mini-batch of data with an equal composition of labeled and unlabeled scenes at each iteration.
 The unlabeled scenes are assigned with pseudo-labels by the teacher model and then train the student model jointly with labeled scenes.
 We use three types of loss functions to supervise the model training here, which is shown as follows:
-$\ \mathcal{L}_{s}=\mathcal{L}_{seg}(\Phi_{\theta_{s}}(x_{i}^{l}),y_{i}^{l}),$ (7)
-$\ \mathcal{L}_{l}=\mathcal{L}_{seg}(\Phi_{\theta_{s}}(x_{i}^{lpf}),y_{i}^{lpf}),$ (8)
-$\ \mathcal{L}_{u}=\mathcal{L}_{seg}(\Phi_{\theta_{s}}(\hat{x}_{i}^{upf}),\hat{y}_{i}^{upf}),$ (9)
+$\ \mathcal{L}_{s}=\mathcal{L}_{seg}(\Phi_{\theta_{\mathbbm{s}}}(x_{i}^{l}),y_{i}^{l}),$ (7)
+$\ \mathcal{L}_{l}=\mathcal{L}_{seg}(\Phi_{\theta_{\mathbbm{s}}}(x_{i}^{lpf}),y_{i}^{lpf}),$ (8)
+$\ \mathcal{L}_{u}=\mathcal{L}_{seg}(\Phi_{\theta_{\mathbbm{s}}}(\hat{x}_{i}^{upf}),\hat{y}_{i}^{upf}),$ (9)
 where $\mathcal{L}_{s}$, $\mathcal{L}_{l}$, and $\mathcal{L}_{u}$ denote the segmentation loss for the labeled scene $x_{i}^{l}$, augmented labeled scene $x_{i}^{lpf}$, and augmented unlabeled scene $\hat{x}_{i}^{upf}$, respectively, while $\mathcal{L}_{seg}$ is the original loss of the segmentation network $\Phi(\cdot)$.
-We sum up all the losses together as our final training target that
+We sum up all the loss term to define the final training objective as:
 $\mathcal{L}=\mathcal{L}_{s}+\lambda_{u}\mathcal{L}_{u}+\lambda_{l}\mathcal{L}_{l},$ (10)
-where $\lambda_{u}$ and $\lambda_{l}$ are loss weights that balance the two loss terms. After that, we update the teacher model weights via the EMA strategy described in Eq. (1).
-SemanticKITTI nuScenes
-Method Venue 1% 5% 10% 20% 40% 50% 1% 10% 20% 50%
-Cylinder3D CVPR 2021 45.4 45.4 56.1 57.8 58.7 58.7 50.9 65.9 66.6 71.2
-MeanTeacher NeurIPS 2017 45.4 - 57.1 59.2 - 60.0 51.6 66.0 67.1 71.7
-CBST ECCV 2018 48.8 - 58.3 59.4 - 59.7 53.0 66.5 69.6 71.6
-CPS CVPR 2021 46.7 - 58.7 59.6 - 60.5 52.9 66.3 70.0 72.5
-GPC ICCV 2021 - 41.8 49.9 58.8 59.9 - - - - -
-ScribbleKITTI CVPR 2022 - 49.9 58.7 59.1 60.9 - - - - -
-LiM3D CVPR 2023 58.4 59.5 62.2 63.1 63.3 63.6 - - - -
-CylinderMix* AAAI 2024 51.3 56.5 60.8 62.0 62.7 62.5 55.7 69.5 72.0 73.3
-LaserMix CVPR 2023 50.6 56.7 60.0 61.9 62.3 62.3 55.3 69.9 71.8 73.2
-Ours (CoScene) - 54.5 62.1 63.3 63.7 63.8 64.3 56.6 70.2 72.8 73.9
-Table 1: Experimental results on the SemanticKITTI and nuScenes dataset compared with different SSL methods.
-The results are reported with mIoU scores (%) and the best score for different split is marked in bold.
-“*” denotes that we reproduce the CylinderMix in the LiDAR branch (Chen et al.
-2024) for fair comparison.
-Experiments
-Datasets and Evaluation Metrics
+where $\lambda_{u}$ and $\lambda_{l}$ are loss weights that balance the two loss terms. After that, we update the teacher model weights using the EMA strategy described in Eq. 1.
+4 Experiments
+4.1 Datasets and Evaluation Metrics
 SemanticKITTI.
-Following state-of-the-art methods (Kong et al.
-2023b; Chen et al.
-2024), we evaluate our CoScene on the SemanticKITTI dataset (Behley et al.
-2019), which is a large-scale segmentation dataset consisting of LiDAR data derived from the popular KITTI dataset (Geiger, Lenz, and Urtasun 2012).
+Following state-of-the-art methods [15, 21], we evaluate our AIScene on the SemanticKITTI dataset [2], which is a large-scale segmentation dataset consisting of LiDAR data derived from the popular KITTI dataset [11].
 It is composed of 22 sequences, with sequences 00-10 serving as the train set (of which 08 as the val set) and sequences 11-21 as the test set.
@@
 nuScenes.
-The nuScenes dataset (Fong et al.
-2022) is another well-known large autonomous driving dataset.
+The nuScenes dataset [10] is another well-known large autonomous driving dataset.
 This dataset is composed of 1000 sequences of 20s duration (850 for training and validation, 150 for testing), which are captured by the LiDAR sensor from Boston and Singapore.
-Following the official splitting protocol, the train and val sets contains 28,130 and 6,019 point cloud scenes, respectively, including per-points annotation for 16 semantic classes.
+Following the official splitting protocol, the train and val sets contain 28,130 and 6,019 point cloud scenes, respectively, including per-points annotation for 16 semantic classes.
 Evaluation Metrics.
-For both the SemanticKITTI and nuScenes datasets, we sample different ratios of annotated train scenes to obtain semi-supervised data and ensure the sampling indices are the same as in LaserMix(Kong et al.
-2023b) for fair comparison.
-Meanwhile, we report the performance on the validation set using the mean Intersection over Union (mIoU) metric as the previous works (Li, Shum, and Breckon 2023; Zhu et al.
-2021).
-Implementation Details
-We conduct experiments with 4$\times$ NVIDIA V100 GPUs with 32GB RAM, based on MMDetection3D (Contributors 2020).
-Inspired by the GT-sampling strategy (Yan, Mao, and Li 2018), we set $\tau_{min}$ to 5 in our experiments.
+For both the SemanticKITTI and nuScenes datasets, we sample different ratios of annotated train scenes to obtain semi-supervised data.
+Meanwhile, we report the performance on the validation set using the mean Intersection over Union (mIoU) metric [21, 23].
+4.2 Implementation Details
+We conduct experiments with 4$\times$ NVIDIA V100 GPUs with 32GB RAM, based on MMDetection3D [8].
+Inspired by the GT-sampling strategy [47], we set $\tau_{min}$ to 5 in our experiments.
 For the threshold $\tau_{s}$ for pseudo-label generation and point erasure, we use the same parameter setting as in LaserMix.
+We use the MinkowskiNet [7] and Cylinder3D [55] as the backbone network due to its widely adopted in previous methods [15, 27, 41].
 To facilitate patch-based data augmentation, we restrict our focus to point clouds within the X and Y dimensions in the range of $[-50,50]m$ for both the SemanticKITTI and nuScenes datasets.
-Similar to mainstream work, we also adopt the consistency loss proposed in MeanTeacher (Tarvainen and Valpola 2017) to align the predictions of the teacher and student networks.
+Similar to mainstream work, we also adopt the consistency loss proposed in MeanTeacher [38] to align the predictions of the teacher and student networks.
 Additionally, the weights of the consistency loss are assigned as $1\times 10^{3}$ and $2\times 10^{3}$ for the SemanticKITTI and nuScenes datasets, respectively.
 The weights for $\lambda_{u}$ and $\lambda_{l}$ are set as 1 for both datasets.
-Main Results
-In this subsection, we compare our method with current SOTA semi-supervised point cloud segmentation methods, using sub-sampling of the datasets under various partition protocols (Kong et al.
-2023b).
-Specifically, in the SemanticKITTI dataset (Behley et al.
-2019), we split the entire training set into labeled subsets with ratios of 1%, 5%, 10%, 20%, 40%, and 50%.
-Similarly, in the nuScenes dataset (Fong et al.
-2022), we utilize labeled subsets with ratios of 1%, 10%, 20%, and 50%.
-We provide detailed comparative results on both datasets against different baseline methods.
-Results on SemanticKITTI Dataset.
-We present the comparison results between our method and state-of-the-art methods on the widely adopted SemanticKITTI dataset in Table 1.
-It can be clearly seen that our method effectively exploits unlabeled scenes compared to the fully supervised learning baseline Cylinder3D (Zhu et al.
-2021), with a dramatic performance boost.
-Specifically, in the 5% ratio, our approach surpasses the fully supervised baseline by 16.7%.
-In other settings, our method also shows consistent improvements ranging from 5.6% to 7.2% for ratios of 10%, 20%, and 50%.
-In addition, comparison to the recently proposed SOTA method LaserMix (Kong et al.
-2023b), which is considered a strong SSL-based point cloud segmentation method, our approach achieves improvements of 2% to 6% in all cases.
-The results also demonstrate that our approach outperforms the current SOTA LiM3D (Li, Shum, and Breckon 2023) in the majority of settings.
-In some partition protocols, our approach shows superior performance compared to the LiM3D with fewer labeled scenes.
-For example, our approach trained with 10% labeled scenes outperforms LiM3D trained with 40% labeled scenes.
-It is worth mentioning that our approach utilizes uniform sampling, similar to LaserMix, to acquire labeled scenes, without utilizing the keyframe sampling method used in LiM3D (Li, Shum, and Breckon 2023), which typically yields higher gains when the amount of labeled data is relatively low.
-Results on nuScenes Dataset.
-As shown in Table 1, our approach still shows a significant improvement in performance compared to state-of-the-art methods on the comprehensive nuScenes dataset (Fong et al.
-2022).
-For instance, our approach outperforms LaserMix (Kong et al.
-2023b) by 1.3% and 1.0% at 1% and 20% ratios, respectively, demonstrating the generalization ability of our method.
-Ablation Study
-In this subsection, we study the effects of intra-scene consistency assurance through the point erasure strategy (PtErase), as well as the exploration of inter-scene correlation, incorporating patch-based mix (MixPatch) and patch-based instance filling data augmentation (InsFill) in our method.
-Without loss of generalizability, all experiments are conducted on the SemanticKITTI dataset with 1%, 10%, 20% and 50% ratio, and evaluated on val set.
-Table 2 demonstrates the improvements achieved by each of these components.
-Ours CoScene Splits on SemanticKITTI Avg.
+Our method excludes all forms of model ensemble and test-time augmentation during the evaluation.
+4.3 Main Results
+In this subsection, we compare our method with current SOTA semi-supervised point cloud segmentation methods (e.g., LaserMix [15], LiM3D [23] and DDSemi [21]), using sub-sampling of the datasets under various partition protocols [21].
+Specifically, we split the entire training set into labeled subsets with ratios of 1%, 10%, 20%, and 50% in the SemanticKITTI [2] and nuScenes dataset [10].
+For further comparison, we also include several classical semi-supervised learning methods [38, 56, 4] from the 2D image domain adapted to 3D segmentation.
+Tab.
+1 summarizes the quantitative results on SemanticKITTI and nuScenes datasets.
+Our method consistently outperforms other comparative methods on both datasets, achieving average mIoU improvements of 1.3% and 2.4%, respectively.
+Notably, the proposed method demonstrates even greater gains on smaller sampling splits, such as 1% and 10%, which are particularly beneficial for the semi-supervised setting.
+Specifically, with a 1% ratio, our method shows mIoU improvements of up to 1.9% and 5.3% on the two datasets, respectively, compared to the recently proposed SOTA method, DDSemi [21].
+These results also validate that our method is less sensitive to reduced labeled data sampling than other methods.
+Furthermore, we present qualitative results in Fig.
+5.
+As indicated by the purple and red dashed circles, our method achieves better segmentation of boundary regions and small objects.
+4.4 Ablation Study
+In this subsection, We first validate the adaptability of our method to different segmentation models.
+We then examine the impact of intra-scene consistency through the point erasure strategy and explore inter-scene correlation by incorporating patch-based mixing and patch-based instance filling data augmentation.
+Figure 5: Qualitative comparison results from LiDAR bird’s eye view on SemanticKITTI val set.
+We set the point with correct and incorrect predicted semantic class in blue and orange for better visualization.
+Best viewed in color.
+Adaptability with Different Segmentation Methods.
+Our AIScene is segmentor-agnostic and can be easily used to further improve off-the-shelf 3D fully-supervised segmentors under the semi-supervised setting.
+Tab.
+2 shows that the proposed method can substantially boost the performance with Cylinder3D [55] as the backbone, indicating the generalization ability of our framework.
+For consistency in comparison, the subsequent experiments are based on Cylinder3D unless otherwise specified.
+Effectiveness of Point Erasure Strategy.
+The 1st row in Tab.
+3 reports the performance of the supervision-only Cylinder3D model [55], which serves as the baseline.
+When PtErase is applied to the baseline, it effectively boosts performance by ensuring consistency during forward and backward propagation, thus addressing the limitations of the pseudo-label mechanism.
+This consistency is achieved through a plug-and-play point erasure strategy, which can be easily integrated into methods relying on the pseudo-label mechanism.
+To further validate the importance of maintaining consistency during forward and backward propagation, we embedded the PtErase strategy into LaserMix [15].
+As illustrated in Fig.
+6 (a), using the PtErase strategy improves performance across all ratios, underscoring the significance of intra-scene consistency and demonstrating the generalizability of our method.
+Moreover, Fig.
+6 (b) shows that the proportion of points removed by the erasure strategy is initially high at the start of training, but gradually decreases as the network converges.
+Furthermore, We present a comparison of class probability distributions in Fig.
+7 to validate the effectiveness of the point erasure strategy.
+The results demonstrate that this strategy reduces classification ambiguity, enabling the model to better capture the features of the correct class.
+Datasets Methods 1% 10% 20% 50% Avg.
+Cylinder3D [55] 45.4 56.1 57.8 58.7 54.5
+Ours (AIScene) 54.5 63.3 63.7 64.3 61.5
+SemanticKITTI Improvement +9.1 +7.2 +5.9 +5.6 +7.0
+Cylinder3D [55] 50.9 65.9 66.6 71.2 63.7
+Ours (AIScene) 56.6 70.2 72.8 73.9 68.4
+nuScenes Improvement +5.7 +4.3 +6.2 +2.7 +4.7
+Table 2: Ablation study on the adaptability of our AIScene using Cylinder3D as the segmentation model.
+Ours AIScene Splits on SemanticKITTI Avg.
 PtErase MixPatch InsFill 1% 10% 20% 50%
-45.4 57.1 59.2 60.0 55.4
+45.4 56.1 57.8 58.7 54.5
 51.7 59.8 61.8 60.4 58.4
@@
 54.5 63.3 63.7 64.3 61.5
-Table 2: Ablation study of different components of our designed CoScene.
-PtErase refers to the point erause strategy, while MixPatch and InsFill refers to the patch mix and instance filling data augmentation, respectively.
-All mIoU results are reported in percentage (%).
-Figure 5: Performance comparison on SemanticKITTI val split under different semi-supervised settings.
-(a) Illustration of performance differences whether using consistency assurance strategy on the LaserMix (Kong et al.
-2023b).
-(b) Comparison between dual-scene level and multi-scene level for the patch-based data augmentation in our CoScene network.
-(c) Results of mIoU against the number of patch split $n$ under the 10% SemanticKITTI split.
-Effectiveness of point erasure strategy.
-The row 1 in Table 2 denotes the performance of MeanTeacher (Tarvainen and Valpola 2017) as the baseline, revealing a marginal enhancement compared to the supervision-only Cylinder3D model (Zhu et al.
-2021).
-When PtErase is applied to the baseline, it is effective and can boost the performance by ensuring consistency during forward and backward propagation, thereby compensating for the shortcomings of the pseudo-label mechanism.
-The consistency assurance is implemented through a plug-and-play point erasure strategy, which can be easily integrated into methods based on the pseudo-label mechanism.
-To further validate the significance of ensuring consistency during forward and backward propagation, we embedded the PtErase strategy into LaserMix.
-As illustrated in Figure 5 (a), performance improves across all ratios when using the PtErase strategy.
-This further emphasizes the importance of intra-scene consistency and demonstrates the generalization ability of our method.
+Table 3: Ablation study of different components of our AIScene framework.
+PtErase refers to the point erasure strategy, while MixPatch and InsFill refer to the patch mix and instance filling data augmentation, respectively.
+[55]
+[49]
+[48]
+[26]
+[45]
+[15]
+[5]
+Figure 6: (a) Illustration of performance differences whether using point erasure strategy on the LaserMix [15].
+(b) Variation in the proportion of points removed by the point erasure strategy throughout the training process.
+(c) Comparison of different mixing-based techniques.
+(d) Results of mIoU against the number of patch split $n$ under the 10% SemanticKITTI split.
+Figure 7: The left chart shows the initial output with distributed probabilities across multiple classes, indicating uncertainty in class prediction.
+Through our point erasure strategy, the right chart exhibits a distribution with a dominant class (“person”) having a significantly higher probability,
 Effectiveness of Patch-based Data Augmentation.
-The patch-based data augmentation consists of MixPatch (patch-based mix) and InsFill (instance filling).
-Rows 3 and 4 of Table 2 demonstrate that using MixPatch or InsFill alone results in an average mIoU improvement of 3.1% or 3.4%, respectively, compared to the baseline.
-This suggests that annotation information transfer across scenes benefits SSL 3D segmentation.
-As shown in rows 5 and 6 of the Table 2, combining MixPatch or InsFill with PtErase significantly enhances performance, which indicates the effectiveness of the patch-based data augmentation.
-Furthermore, the collaboration of all three components, leads to substantial performance gains, as validated by the last row of Table 2.
-Additionally, our patch pool and instance pool within the patch-based data augmentation facilitate effective information fusion across multiple scenes.
-This approach inherently enables the transfer of inter-scene diverse correlation information.
-To verify the effectiveness of the interaction at the multi-scene level, we modify the pool implementation to enable only interaction between two scenes.
-The comparative results, as shown in Figure 5 (b), indicate superior performance at multi-scene level compared to dual-scene level.
+The patch-based data augmentation consists of MixPatch (patch-based mixing) and InsFill (instance filling).
+In Tab.
+3, the results in the 3rd and 4th rows indicate that using MixPatch or InsFill individually yields an average mIoU improvement of 4.0% or 4.3%, respectively, over the baseline.
+This suggests that incorporating diverse semantic information across scenes benefits semi-supervised LiDAR segmentation.
+Furthermore, as shown in the 5th and 6th rows, combining MixPatch or InsFill with PtErase significantly enhances performance, highlighting the effectiveness of patch-based data augmentation.
+The combination of all three components yields substantial performance gains, as evidenced by the last row.
+Additionally, the patch pool and instance pool within the patch-based data augmentation facilitate the fusion of diverse semantic information across multiple scenes.
+To verify the effectiveness of interactions at the multi-scene level, we modify the pool implementation to allow interactions between only two scenes.
+The results in Tab.
+4 indicate that multi-scene interactions yield superior performance compared to dual-scene interactions.
 Different Mixing Strategies.
-Figure 5 (c) compares the patch-based data augmentation employed in our CoScene with other mixing approaches (Nekrasov et al.
-2021; Xiao et al.
-2022; Yun et al.
-2019; Zhang et al.
-2017).
-MixUp demonstrates a slight improvement over the baseline, likely due to its limited ability to transform information.
-Other methods have a considerable improvement over the baseline by enriching point cloud distributions.
-In contrast, our CoScene achieves the greatest gain by effectively exploring rich inter-scene correlation, highlighting its effectiveness in facilitating information transfer between labeled and unlabeled scenes.
-This is accomplished through its innovative design, which includes a patch pool and an instance pool spanning multiple scenes.
+Fig.
+6 (c) compares our patch-based data augmentation with existing mixing methods [26, 45, 48].
+MixUp [49] shows a slight improvement over the baseline, likely due to its limited capacity for information transformation.
+Note that we reproduce CylinderMix in the LiDAR branch [5] for comparison.
+Other methods demonstrate considerable improvement over the baseline by enriching point cloud distributions.
+In contrast, our AIScene achieves the most substantial gains by effectively leveraging inter-scene correlation, highlighting its effectiveness in facilitating diverse semantic information generation for labeled and unlabeled scenes.
+This is achieved through an innovative design that incorporates a patch pool and an instance pool spanning multiple scenes.
+Splits on Semantic KITTI
+Augmentation Type Methods 1% 10% 20% 50% Avg.
+dual-scene level Ours 53.4 61.3 61.8 62.2 59.7
+multi-scene level Ours 54.5 63.3 63.7 64.3 61.5
+Improvement - +1.1 +2.0 +1.9 +2.1 +1.8
+Table 4: Comparison between dual-scene and multi-scene levels in the patch-based data augmentation of our AIScene.
 Patch Split Number.
-As the patch split number $n$ plays a critical role in our framework, we further analyze the impact of varying $n$ in patch-based data augmentation, with the results presented in Figure 6.
-When $n$ is small, it can yield a certain level of enhancement, but it may fail to generate data sufficiently divergent from the original, impeding the effective transfer of information between labeled and unlabeled scenes.
-Conversely, when $n$ is too large, the generated scenes may deviate excessively from the original data, resulting in the loss of essential prior scene information.
-Empirically, we identify that setting $n$ to 18 yields optimal results.
-Therefore, throughout this paper, we consistently use $n$ set to 18, unless explicitly stated otherwise.
-Figure 6: Comparison of different mixing-based techniques.
-Conclusion
-In this paper, we explore the scene coherence for semi-supervised point cloud segmentation, building upon the intra-scene consistency and inter-scene correlation.
-In particular, we achieve intra-scene consistency assurance through a point erasure strategy that ensures consistency during the forward and backward propagation, rather than simply disregarding points with pseudo-labels as in previous methods.
-Additionally, we explore inter-scene correlation and propose a patch-based data augmentation to enhance information transfer between labeled scenes and unlabeled scenes.
-Benefiting from these two components, our CoScene effectively transfers diverse semantic information and produces more accurate segmentation results.
-Extensive experimental results demonstrate the superiority of our CoScene over state-of-the-art methods.
+As the patch split number $n$ plays a critical role in our framework, we further analyze the impact of varying $n$ in patch-based data augmentation, with the results presented in Fig.
+6 (d).
+When $n$ is small, it can yield a certain level of enhancement, but it may fail to generate data sufficiently divergent from the original, thereby impeding the effective generation of diverse semantic information between labeled and unlabeled scenes.
+Conversely, when $n$ is too large, the generated scenes may deviate excessively from the original data, resulting in a loss of essential prior scene information.
+Empirically, we find that setting $n$ to 18 yields optimal results.
+Throughout this paper, we consistently use $n$ set to 18 unless explicitly stated otherwise.
+5 Conclusion
+This paper explores scene affinity for semi-supervised LiDAR semantic segmentation in driving scenes, named AIScene.
+Specifically, AIScene addresses intra-scene inconsistency through a point erasure strategy, which enhances the learning of pseudo-labeled points.
+Moreover, it employs a patch-based data augmentation method to mix multiple scenes at both scene and instance levels, guided by inter-scene correlation.
+This method enriches the semantic diversity of labeled scenes and complements pseudo-labeled scenes.
+Extensive experiments show that AIScene achieves state-of-the-art results on popular benchmarks, with remarkable improvements in the most challenging settings.
+Additionally, the novel erasure strategy and augmentation method can function independently to facilitate semi-supervised LiDAR semantic segmentation.

### v3

@@
 Compared to existing augmentation techniques that typically perform scene-level mixing between two scenes, our method enhances the semantic diversity of labeled (or pseudo-labeled) scenes, thereby improving the semi-supervised performance of segmentation models.
-Experiments show that AIScene outperforms previous methods on two popular benchmarks across four settings, achieving notable improvements of 1.9% and 5.3% in the most challenging 1% labeled data.
-1 Introduction
-Semantic segmentation of LiDAR point clouds is essential for autonomous driving, allowing vehicles to perceive their 3D surroundings.
-Most state-of-the-art models for driving-scene LiDAR segmentation are dependent on high-quality, point-annotated datasets [12, 55, 35].
-With numerous objects in driving scenes, densely annotating points is extremely labor-intensive.
-Consequently, semi-supervised LiDAR semantic segmentation has attracted growing attention, where models are trained on a limited amount of labeled data supplemented by a large set of unlabeled data.
+Experiments show that AIScene outperforms previous methods on two popular benchmarks across four settings, achieving notable improvements of 1.9% and 2.1% in the most challenging 1% labeled data.
+The code will be released at https://github.com/azhuantou/AIScene.
+footnotetext: Equal contribution.  Corresponding authors.
 Figure 1: Top: Pseudo-labeled scenes generated from unlabeled scenes typically contain points with pseudo labels (colored) and without them (gray).
@@
 Their resulting scenes may be constrained in terms of semantic diversity, potentially lacking objects like bicyclists and buses, which are important in driving scenes.
+1 Introduction
+Semantic segmentation of LiDAR point clouds is essential for autonomous driving, allowing vehicles to perceive their 3D surroundings.
+Most state-of-the-art models for driving-scene LiDAR segmentation are dependent on high-quality, point-annotated datasets [12, 57, 36].
+With numerous objects in driving scenes, densely annotating points is extremely labor-intensive.
+Consequently, semi-supervised LiDAR semantic segmentation has attracted growing attention, where models are trained on a limited amount of labeled data supplemented by a large set of unlabeled data.
 The common solution for learning from unlabeled data is to apply teacher-student training, where a teacher network generates pseudo-labeled scenes from unlabeled data, which are then used to supervise the learning of a student network.
@@
 AIScene has been extensively evaluated through experiments and demonstrated superiority on two popular benchmarks for semi-supervised LiDAR semantic segmentation.
-It exceeds state-of-the-art methods by 1.3% on the SemanticKITTI [2] and 2.4% on the nuScenes [10] in terms of average performance across four settings (1%, 10%, 20%, and 50% labeled data).
-Notably, in the most challenging setting, with only 1% labeled data, AIScene achieves remarkable improvements of 1.9% and 5.3% on the two datasets.
+It exceeds state-of-the-art methods by 1.3% on the SemanticKITTI [2] and 1.2% on the nuScenes [10] in terms of average performance across four settings (1%, 10%, 20%, and 50% labeled data).
+Notably, in the most challenging setting, with only 1% labeled data, AIScene achieves remarkable improvements of 1.9% and 2.1% on the two datasets.
 Additionally, our ablation studies confirm AIScene’s potential to reduce the requirements for labeled data across different LiDAR semantic segmentation networks (see Tab.
@@
 To address this, our AIScene leverages a point erasure strategy to mitigate this inconsistency, complementing the pseudo-label mechanism.
-3D Point Cloud Data Augmentation: Point cloud data augmentation has been explored across various 3D vision tasks [54, 33, 42, 9].
+3D Point Cloud Data Augmentation: Point cloud data augmentation has been explored across various 3D vision tasks [56, 34, 44, 9, 43, 25].
 Inspired by the success of Mixup [51] in image classification, several methods have investigated techniques for mixing selected point cloud samples.
@@
 This prevents low-quality pseudo labels from persisting in pools and misleading model training.
-Additionally, if the number of points within a patch or instance falls below a certain threshold $\tau_{\text{min}}$, it will not be saved into the pool, ensuring effective semantic diversity.
+Additionally, if the number of points within a patch or instance falls below a certain threshold $\tau_{\text{min}}$, it will not be saved into the pool.
+In practice, we set $\tau_{\text{min}}$ to 5 [49].
+If a patch or instance has fewer than 5 points, its spatial structure exhibits high uncertainty, making it semantically ambiguous.
+Thus, it is excluded from the pools to prevent noise during data augmentation.
 Figure 4: (a) Illustration of the MixPatch process.
@@
 DDSemi [21] CVPR’2024 59.3 65.1 66.3 67.0 64.4 58.1 70.2 74.0 76.5 69.7
-Ours (AIScene) - 61.2 66.3 67.4 67.9 65.7 63.4 73.7 75.0 76.3 72.1
-Improvement $\uparrow$ - +1.9 +1.2 +1.1 +0.9 +1.3 +5.3 +3.5 +1.0 -0.2 +2.4
+Ours (AIScene) - 61.2 66.3 67.4 67.9 65.7 60.2 72.3 75.0 76.4 70.9
+Improvement $\uparrow$ - +1.9 +1.2 +1.1 +0.9 +1.3 +2.1 +2.1 +1.0 -0.1 +1.2
 Table 1: Experimental results on the SemanticKITTI and nuScenes dataset compared with different SSL methods.
 The results are reported with mIoU scores (%).
-The blod and underline scores respectively denote the best and second results.
+The bold and underline scores respectively denote the best and second results.
 3.4 Model Training
@@
 We conduct experiments with 4$\times$ NVIDIA V100 GPUs with 32GB RAM, based on MMDetection3D [8].
-Inspired by the GT-sampling strategy [47], we set $\tau_{min}$ to 5 in our experiments.
-For the threshold $\tau_{s}$ for pseudo-label generation and point erasure, we use the same parameter setting as in LaserMix.
+For the threshold $\tau_{s}$ for pseudo-label generation and point erasure, we experimentally find that 0.9 yields relatively satisfactory results.
 We use the MinkowskiNet [7] and Cylinder3D [57] as the backbone network due to its widely adopted in previous methods [15, 28, 42].
@@
 1 summarizes the quantitative results on SemanticKITTI and nuScenes datasets.
-Our method consistently outperforms other comparative methods on both datasets, achieving average mIoU improvements of 1.3% and 2.4%, respectively.
+Our method consistently outperforms other comparative methods on both datasets, achieving average mIoU improvements of 1.3% and 1.2%, respectively.
 Notably, the proposed method demonstrates even greater gains on smaller sampling splits, such as 1% and 10%, which are particularly beneficial for the semi-supervised setting.
-Specifically, with a 1% ratio, our method shows mIoU improvements of up to 1.9% and 5.3% on the two datasets, respectively, compared to the recently proposed SOTA method, DDSemi [21].
+Specifically, with a 1% ratio, our method shows mIoU improvements of up to 1.9% and 2.1% on the two datasets, respectively, compared to the recently proposed SOTA method, DDSemi [21].
 These results also validate that our method is less sensitive to reduced labeled data sampling than other methods.
@@
 (b) Variation in the proportion of points removed by the point erasure strategy throughout the training process.
-(c) Comparison of different mixing-based techniques.
+(c) Comparison of different mixing-based techniques on SemanticKITTI dataset.
 (d) Results of mIoU against the number of patch split $n$ under the 10% SemanticKITTI split.
@@
 This method enriches the semantic diversity of labeled scenes and complements pseudo-labeled scenes.
-Extensive experiments show that AIScene achieves state-of-the-art results on popular benchmarks, with remarkable improvements in the most challenging settings.
+Extensive experiments show that AIScene achieves state-of-the-art results on popular benchmarks.
 Additionally, the novel erasure strategy and augmentation method can function independently to facilitate semi-supervised LiDAR semantic segmentation.
+Acknowledgments. This work is supported by the National Natural Science Foundation of China (No.62325111, U22B2011, 62271354).